find the derivative using chain rule.



1. y= (x³ + 3)⁵



2. y=(-3x⁵ + 1)³



3. y=(-5x³ - 3)³



4. y=(5x² + 3)⁴



5. f(x)=⁴√-3x⁴-2



6. f(x)= √-2x² + 1




1
Expert's answer
2022-04-27T11:29:17-0400

"1. y=(x^3+3)^5\\
y'=5(x^3+3)^4(x^3+3)' = 15x^2(x^3+3)^4"

"2. y =(-3x^5+1)^3\\
y'=3(-3x^5+1)^2(-3x^5+1)' = -45x^4(-3x^5+1)^2"

"3. y =(-5x^3-3)^3\\
y'=3(-5x^3-3)^2(-5x^3-3)'= -45x^2(-5x^3-3)^2"

"4. y=(5x^2+3)^4\\
y'=4(5x^2+3)^3(5x^2+3)' = 40x(5x^2+3)^3"

"5. f(x) = \sqrt[4]{-3x^4-2}\\
f'(x)=\cfrac{(-3x^4-2)'}{4\sqrt[4]{(-3x^4-2)^3}}=\cfrac{-3x^3}{\sqrt[4]{(-3x^4-2)^3}}"

"6. f(x) = \sqrt{-2x^2+1}\\
f'(x) = \cfrac{(-2x^2+1)'}{2\sqrt{-2x^2+1}} = \cfrac{-2x}{\sqrt{-2x^2+1}}"